OpenStudy (anonymous):

Convert the complex number 2 - 2 sqrt3 i into its polar representation.

OpenStudy (anonymous):

4(cos(300) + isin(300))

OpenStudy (earthcitizen):

what are you doing ?

OpenStudy (earthcitizen):

\[r<\theta \]

OpenStudy (earthcitizen):

\[ \theta =\tan^{-1} (-\sqrt{3})=-60\] \[r=4\]

OpenStudy (earthcitizen):

\[ \therefore 2-i2\sqrt{3} = 4<-60\]